True or False:

 Consider the integral "\displaystyle{\int \frac{x^3+x}{x-1}}dx" To evaluate the integral, we need to first perform the long division.


1
Expert's answer
2021-09-08T10:14:31-0400

True


"\frac{x^3+x}{x-1}=x^2+\frac{x^2+x}{x-1}\\
=x^2+x+2+\frac{2}{x-1}\\
\frac{x^3+x}{x-1}=x^2-x+2-\frac{2}{x+1}\\
\mathrm{Apply\:the\:Sum\:Rule}:\quad \int f\left(x\right)\pm g\left(x\right)dx=\int f\left(x\right)dx\pm \int g\left(x\right)dx\\
=\frac{x^3}{3}-\frac{x^2}{2}+2x-2\ln \left|x+1\right|"